Find the Enthalpies of Combustion of gasoline (C8H18), methane (CH4), methnol (CH3OH), ethanol (C2H5OH), propane (C3H8) and kerosene (C15H32). Then determine the Energy-to-Pollution Ratios (EPRs) of each fuel and compare them. Discuss which fuels are most rational to use...if we insist upon burnign hydrocarbons to obtain energy.


1
Expert's answer
2021-02-03T07:16:08-0500

Enthalpy of combustion is calculated using the following formula,

"\Delta H^0_{rxn}=\sum\Delta H^0_f(products)-\sum\Delta H^0_f(reactants)"

Gasoline

"2C_8H_{18(l)}+25O_{2(g)}\to16CO_{2(g)}+18H_2O_{(l)}\newline \Delta H^0_c=4730kJ/g"

Methane

"CH_{4g}+O_{2g}\to CO_{2g}+2H_2O_l+\Delta H^0_{rxn}\newline =[(-393.5-2*286)-(-74.85)\newline= 890.7kJ/g"

Methanol

"CH_3OH_g+\dfrac{3}{2}O_{2g}\to CO_{2g}+2H_2O_l\newline \newline =726kJ/g"

Ethanol

"C_2H_5OH_g+3O_{2g}\to 2CO_{2g}+3H_2O_l\newline =1368kJ/g"

Propane

"C_3H_{8g}+5O_{2g} \to 3CO_{2g} +4H_2O_l\newline =2220kJ/g"